# Local uniqueness and non-degeneracy of blowup solutions for regular Liouville systems

### v1

@@ 3.5 A refined expansion of $\Lambda_{I,N}(\rho^{k})$
$\Lambda_{I,N}(\rho^{k})$ $=\sum_{i=1}^{n}\sum_{t=1}^{N}\frac{2-m^{*}_{i}}{N}D_{it}^{k}\frac{e^{H_{i,t}(q^{M,k})}}{e^{H_{i,1}(q^{M,k})}}(\varepsilon_{1}^{k})^{m_{i}^{k}-2}+T^{\Lambda}_{k,M}+O(\varepsilon^{M(m^{*}-2){-\delta}}).$
@@
and $T_{k,m}^{\Lambda}=o_{\tau}(1)\varepsilon^{m^{*}-2}$ and is $C^{2}$ depends on $p_{i}^{M,k}$, $h_{i}$ and its derivatives up to $M+3$ order. If $m^{*}=4$, we get
$\Lambda_{I,N}(\rho^{k})=-\frac{2}{N}\sum_{i=1}^{n}\sum_{t=1}^{N}\Bigg{[}b^{k}_{it}(\varepsilon_{t}^{k})+D_{it}^{k}\frac{e^{H_{it}(q^{M,k})}}{e^{H_{i1}(q^{M,k})}}(\varepsilon_{1}^{k})^{2}\Bigg{]}+T^{\Lambda}_{k,M}+O(\varepsilon^{2M-\delta}).$ (3.50)
@@
Here, $T_{i,M}=O(\varepsilon^{2m^{*}-4})$ and depends $C^{2}$-continuously on $p_{t}^{M,k}$ and the derivatives of $h_{i}^{k}$ up to $M+3$-th order. As for the integrals $\rho_{ti}^{k}$’s, by Theorem 3.1, we get
$\rho_{it}^{k}$ $=\rho_{i}^{k}\int_{B_{\tau_{k}^{M}}^{M}(q_{t}^{M,k})}h_{i}^{k}e^{u_{i}^{k}}dV_{g}=\int_{B_{{\tau_{k}^{M}/\varepsilon_{t}^{k}}}(0)}H_{t,i}^{M,k}(y)e^{\phi_{i,t}^{k}(p_{t}^{k}+\varepsilon_{t}^{k}y)}e^{v_{i}^{k}(y)}dy$
$=H_{t,i}^{M,k}(0)\int_{\mathbb{R}^{2}}e^{V_{t,i}^{k}}-{\frac{2\pi H_{t,i}^{1,k}(0)e^{I_{i}^{k}}}{m_{i}^{k}-2}\Big{(}\frac{\varepsilon_{t}^{k}}{\tau_{k}^{M}}\Big{)}^{m_{i}^{k}-2}}+\alpha_{t,i}^{k}(\varepsilon_{t}^{k};\tau)+O(\varepsilon^{M(m^{*}-2){-\delta}}).$
Here, $H_{t,i}^{k}(\cdot)=\frac{h^{k}_{i}(q_{t}^{M,k}+\cdot)}{h_{i}^{k}(q_{t}^{k})}e^{\psi+f_{i}^{k}}$ and $\alpha_{i}^{k}(\varepsilon_{t}^{k};\tau)=\int_{\partial\Omega_{M,k}}\frac{\partial c_{2,0,k,i}}{\partial{\bf n}}+\int_{\partial\Omega_{M,k}}\frac{\partial c_{*,0,k,i}}{\partial{\bf n}}$. Then, we get
@@
$\quad-\sum_{t=1}^{N}\frac{2\pi H_{i}^{M,k}(0)e^{I_{i}^{k}}}{m_{i}^{k}-2}\Big{(}\frac{\varepsilon_{t}^{k}}{\tau_{k}^{M}}\Big{)}^{m_{i}^{k}-2}+\sum_{t=1}^{N}\alpha_{t,i}^{k}(\varepsilon_{k};\tau)+T_{i,M}+O(\varepsilon^{M(m^{*}-2){-\delta}}).$
@@
$\quad-\frac{4}{N}\sum_{i=1}^{n}\sum_{t=1}^{N}\frac{e^{I_{i}^{k}}}{m_{i}^{k}-2}\Big{(}\frac{\varepsilon_{t}^{k}}{\tau_{k}^{M}}\Big{)}^{m_{i}^{k}-2}+\frac{4}{2\pi N}\sum_{t=1}^{N}\sum_{i=1}^{n}\big{(}\alpha_{t,i}^{k}(\varepsilon_{t}^{k};\tau)+T_{i,M}\big{)}+O(\varepsilon^{M(m^{*}-2){-\delta}}).$
@@
$\quad-\frac{1}{N}\sum_{i=1}^{n}\sum_{t=1}^{N}\frac{2m_{t,i}^{k}}{m_{t,i}^{k}-2}\Big{(}\frac{\varepsilon_{t}^{k}}{\tau}\Big{)}^{m_{i}^{k}-2}+\frac{2}{2\pi N}\sum_{t=1}^{N}\sum_{i=1}^{n}m_{t,i}^{k}\big{(}\alpha_{t,i}^{k}(\varepsilon_{t}^{k};\tau)+T_{i,M}\big{)}+O(\varepsilon^{M(m^{*}-2){-\delta}}).$
@@
$\Lambda_{I,N}(\rho^{k})=4\sum_{i=1}^{n}\frac{\rho_{i}^{k}}{2\pi N}-\sum_{i,j=1}^{n}a_{ij}\frac{\rho_{i}^{k}}{2\pi N}\frac{\rho_{j}^{k}}{2\pi N}=:\sum_{i=1}^{n}\sum_{t=1}^{N}D_{it}^{k}\frac{e^{{H_{i,t}(q^{M,k})}}}{e^{H_{i,1}(q^{M,k})}}(\varepsilon_{1}^{k})^{m_{i}^{k}-2}+T^{\Lambda}_{k,M}+O(\varepsilon^{M(m^{*}-2){-\delta}})$
with $T_{k,m}^{\Lambda}=o_{\tau}(1)\varepsilon^{m^{*}-2}$ and is $C^{2}$ depends on $p_{i}^{M,k}$, $h_{i}$ and its derivatives up to $M+3$ order. Notice that in the last step we use Corollary 3.1. This proves the result.
@@ 3.6 A refined Pohozaev identity
$\int_{B_{\tau_{M}^{k}/\varepsilon_{1}^{k}}(p_{t}^{M,k})}\partial_{1}H_{i}^{M,k}(y)e^{V_{t,i}^{M,k}}+T_{\mbox{Po},M}^{k}+O(\varepsilon^{(M+1)(m^{*}-2)-\delta})=0$ (3.54)
for certain term $T_{\mbox{Po},M}^{k}$ of order $\varepsilon^{m^{*}-2}$ for the case $m^{*}<4$ and $\varepsilon^{2}\ln\frac{1}{\varepsilon}$ for $m^{*}=4$ and depends $C^{1}$-continuously in $p_{t}^{M,k}$, $H^{M,k}_{j}$ and its derivatives up to $M+2$ order.
@@
$\int_{B_{\tau_{M}^{k}/\varepsilon_{1}^{k}}(p_{t}^{M,k})}\partial_{1}H_{i}^{M,k}(y)e^{V_{t,i}^{M,k}}+T_{\mbox{Po},M}^{k}+O(\varepsilon^{(M+1)(m^{*}-2)-\delta})=0$
for certain term $T_{\mbox{Po},M}^{k}$ of order $\varepsilon^{m^{*}-2}$ for the case $m^{*}<4$ and $\varepsilon^{2}\ln\frac{1}{\varepsilon}$ for $m^{*}=4$ and depends $C^{1}$-continuously in $p_{t}^{M,k}$, $H^{M,k}_{j}$ and its derivatives up to $M+2$ order.
@@ 4.1 Comparison between $\varepsilon_{t}^{(1),k}$ and $\varepsilon_{t}^{(2),k}$
$\Lambda_{I,N}(\rho^{k})$ $=\sum_{i=1}^{n}\sum_{t=1}^{N}D_{it}^{k,(1)}\frac{e^{H_{it}(p^{(1),k})}}{e^{H_{i1}(p^{(1),k})}}(\varepsilon_{1}^{(1),k})^{m_{i}^{k}-2}+\sum_{t=1}^{N}\sum_{i=1}^{n}\frac{(2-m_{t,i}^{(1),k})}{\pi N}\big{(}\alpha_{t,i}^{(1),k}(\varepsilon_{t}^{k};\tau)+T^{(1)}_{i,M}\big{)}$
$\quad\quad+O((\varepsilon_{t}^{(2),k})^{M(m^{*}-2){-\delta}})$
$=\sum_{i=1}^{n}\sum_{t=1}^{N}D_{it}^{k,(2)}\frac{e^{H_{it}(p^{(2),k})}}{e^{H_{i1}(p^{(2),k})}}(\varepsilon_{1}^{(2),k})^{m_{i}^{k}-2}+\sum_{t=1}^{N}\sum_{i=1}^{n}\frac{(2-m_{t,i}^{(2),k})}{\pi N}\big{(}\alpha_{t,i}^{(2),k}(\varepsilon_{t}^{k};\tau)+T^{(2)}_{i,M}\big{)}$
$\quad\quad+O((\varepsilon_{t}^{(2),k})^{M(m^{*}-2){-\delta}}).$ (4.1)
Here, the terms $D_{it}^{k,(1)}$, $D_{it}^{k,(2)}$, $\alpha_{t,i}^{(1),k}(\varepsilon_{t}^{k};\tau)$, $T^{(1)}_{i,M}$, $\alpha_{t,i}^{(2),k}(\varepsilon_{t}^{k};\tau)$ and $T^{(2)}_{i,M}$ are defined as in Theorem 3.2.
@@
Notice that the quantities $\alpha_{i}^{(l),k}(\varepsilon_{t}^{(l),k};\tau)$ depends on $p^{(l),k}_{t}$ and $\varepsilon_{t}^{(l),k}$ for $l=1,2$. Then, a direct computation implies that
$|\alpha_{i}^{(1),k}(\varepsilon_{t}^{(1),k};\tau)-\alpha_{i}^{(2),k}(\varepsilon_{t}^{(2),k};\tau)|\leq A+B.$
@@
$|\alpha_{i}^{(1),k}(\varepsilon_{t}^{(1),k};\tau)-\alpha_{i}^{(2),k}(\varepsilon_{t}^{(2),k};\tau)|$ $=O\Bigg{(}\varepsilon^{m^{*}-1}\Bigg{|}p_{t}^{(1),k}-\frac{\varepsilon_{t}^{(2),k}}{\varepsilon_{t}^{(1),k}}p_{t}^{(2),k}\Bigg{|}\Bigg{)}.$
@@ 4.2 Comparison between $p_{t}^{(1),k}$ and $p_{t}^{(2),k}$
$\int_{B_{\tau_{k}^{M}/\varepsilon_{t}^{(1),k}}(p_{t}^{(1),k})}\partial_{1}H_{i}^{M,k}(y)e^{V_{t,i}^{(1),k}}=-T_{\mbox{Po},M}^{(1),k}+O(\varepsilon^{(M+1)(m^{*}-2)-\delta})$
@@
$\int_{B_{\tau_{k}^{M}/\varepsilon_{t}^{(2),k})}(p_{t}^{(2),k})}\partial_{1}H_{i}^{M,k}(y)e^{V_{t,i}^{(2),k}}=-T_{\mbox{Po},M}^{(2),k}+O(\varepsilon^{(M+1)(m^{*}-2)-\delta}).$
@@
$-T_{\mbox{Po},M}^{(1),k}+T_{\mbox{Po},M}^{(2),k}=O(\varepsilon^{m^{*}-2})\Bigg{(}\Bigg{|}1-\frac{\varepsilon_{t}^{(1),k}}{\varepsilon_{t}^{(2),k}}\Bigg{|}+|\varepsilon_{t}^{(1),k}p_{t}^{(1),k}-\varepsilon_{t}^{(2),k}p_{t}^{(2),k}|\Bigg{)}.$ (4.3)
@@ A.3 A remark on a non-vanishing blowup result
In [12], the authors glue a blowup solution under the assumptions $(\mathcal{H}_{1})$, $(\mathcal{H}_{2})$, (2.1), (2.2) and (3.13). The assumption (2.1) does not require

### v2

@@ 3.5 A refined expansion of $\Lambda_{I,N}(\rho^{k})$
$\Lambda_{I,N}(\rho^{k})$ $=\sum_{i=1}^{n}\sum_{t=1}^{N}\frac{2-m^{*}_{i}}{N}D_{it}^{k}\frac{e^{H_{i,t}(q^{M,k})}}{e^{H_{i,1}(q^{M,k})}}(\varepsilon_{1}^{k})^{m_{i}^{k}-2}+T^{\Lambda}_{M,k}+O(\varepsilon^{M(m^{*}-2){-\delta}}).$
@@
and $T_{M,k}^{\Lambda}=o_{\tau}(1)\varepsilon^{m^{*}-2}$ and is $C^{2}$ depends on $p_{i}^{M,k}$, $h_{i}$ and its derivatives up to $M+3$ order. If $m^{*}=4$, we get
$\Lambda_{I,N}(\rho^{k})=-\frac{2}{N}\sum_{i=1}^{n}\sum_{t=1}^{N}\Bigg[b^{k}_{it}(\varepsilon_{t}^{k})+D_{it}^{k}\frac{e^{H_{it}(q^{M,k})}}{e^{H_{i1}(q^{M,k})}}(\varepsilon_{1}^{k})^{2}\Bigg]+T^{\Lambda,*}_{M,k}+O(\varepsilon^{2M-\delta}).$ (3.50)
@@
Here, $T_{M,k,i}^{\rho,out}=O(\varepsilon^{2m^{*}-4})$ and depends $C^{2}$-continuously on $p_{t}^{M,k}$ and the derivatives of $h_{i}^{k}$ up to $M+3$-th order. As for the integrals $\rho_{ti}^{k}$’s, by Theorem 3.1, we get
$\rho_{it}^{k}$ $=\rho_{i}^{k}\int_{B_{\tau_{k}^{M}}^{M}(q_{t}^{M,k})}h_{i}^{k}e^{u_{i}^{k}}dV_{g}=\int_{B_{{\tau_{k}^{M}/\varepsilon_{t}^{k}}}(0)}H_{t,i}^{M,k}(y)e^{\phi_{i,t}^{k}(q_{t}^{M,k}+\varepsilon_{t}^{k}y)}e^{v_{i}^{k}(y)}dy$
$=H_{t,i}^{M,k}(0)\int_{\mathbb{R}^{2}}e^{V_{t,i}^{k}}-{\frac{2\pi H_{t,i}^{1,k}(0)e^{I_{i}^{k}}}{m_{i}^{k}-2}\Big(\frac{\varepsilon_{t}^{k}}{\tau_{k}^{M}}\Big)^{m_{i}^{k}-2}}+T_{M,k,i,t}^{in}+O(\varepsilon^{M(m^{*}-2){-\delta}}).$
Here, $H_{t,i}^{k}(\cdot)=\frac{h^{k}_{i}(q_{t}^{M,k}+\cdot)}{h_{i}^{k}(q_{t}^{M,K})}e^{\psi+f_{i}^{k}}$ and $T_{M,k,i,t}^{in}$ is the term defined similarly as in Theorem 3.1. Then, we get
@@
$\quad-\sum_{t=1}^{N}\frac{2\pi H_{i}^{M,k}(0)e^{I_{i}^{k}}}{m_{i}^{k}-2}\Big(\frac{\varepsilon_{t}^{k}}{\tau_{k}^{M}}\Big)^{m_{i}^{k}-2}+\sum_{t=1}^{N}T_{M,k,i,t}^{in}+T_{M,k,i}^{\rho,out}+O(\varepsilon^{M(m^{*}-2){-\delta}}).$
@@
$\quad-\frac{4}{N}\sum_{i=1}^{n}\sum_{t=1}^{N}\frac{e^{I_{i}^{k}}}{m_{i}^{k}-2}\Big(\frac{\varepsilon_{t}^{k}}{\tau_{k}^{M}}\Big)^{m_{i}^{k}-2}+\frac{4}{2\pi N}\sum_{t=1}^{N}\sum_{i=1}^{n}\big(T_{M,k,i,t}^{in}+T_{M,k,i}^{\rho,out}\big)+O(\varepsilon^{M(m^{*}-2){-\delta}}).$
@@
$\quad-\frac{1}{N}\sum_{i=1}^{n}\sum_{t=1}^{N}\frac{2m_{t,i}^{k}}{m_{t,i}^{k}-2}\Big(\frac{\varepsilon_{t}^{k}}{\tau_{k}^{M}}\Big)^{m_{i}^{k}-2}+\frac{2}{2\pi N}\sum_{t=1}^{N}\sum_{i=1}^{n}m_{t,i}^{k}\big(T_{M,k,i,t}^{in}+T_{M,k,i}^{\rho,out}\big)+O(\varepsilon^{M(m^{*}-2){-\delta}}).$
@@
$\Lambda_{I,N}(\rho^{k})=4\sum_{i=1}^{n}\frac{\rho_{i}^{k}}{2\pi N}-\sum_{i,j=1}^{n}a_{ij}\frac{\rho_{i}^{k}}{2\pi N}\frac{\rho_{j}^{k}}{2\pi N}=:\sum_{i=1}^{n}\sum_{t=1}^{N}D_{it}^{k}\frac{e^{{H_{i,t}(q^{M,k})}}}{e^{H_{i,1}(q^{M,k})}}(\varepsilon_{1}^{k})^{m_{i}^{k}-2}+T^{\Lambda}_{M,k}+O(\varepsilon^{M(m^{*}-2){-\delta}})$
with $T_{M,,k}^{\Lambda}=o_{\tau}(1)\varepsilon^{m^{*}-2}$ and is $C^{2}$ depends on $p_{i}^{M,k}$, $h_{i}$ and its derivatives up to $M+3$ order. Notice that in the last step we use Corollary 3.1. This proves the result.
@@ 3.6 A refined Pohozaev identity
$\int_{B_{\tau_{M}^{k}/\varepsilon_{1}^{k}}(p_{t}^{M,k})}\partial_{1}H_{i}^{M,k}(y)e^{V_{t,i}^{M,k}}+T^{Po}_{M,k}+O(\varepsilon^{(M+1)(m^{*}-2)-\delta})=0$ (3.54)
for certain term $T^{Po}_{M,k}$ of order $\varepsilon^{m^{*}-2}$ for the case $m^{*}<4$ and $\varepsilon^{2}\ln\frac{1}{\varepsilon}$ for $m^{*}=4$ and depends $C^{1}$-continuously in $p_{t}^{M,k}$, $H^{M,k}_{j}$ and its derivatives up to $M+2$ order.
@@
$\int_{B_{\tau_{M}^{k}/\varepsilon_{1}^{k}}(p_{t}^{M,k})}\partial_{1}H_{i}^{M,k}(y)e^{V_{t,i}^{M,k}}+T^{Po}_{M,k}+O(\varepsilon^{(M+1)(m^{*}-2)-\delta})=0$
for certain term $T_{M,k}^{Po}$ of order $\varepsilon^{m^{*}-2}$ for the case $m^{*}<4$ and $\varepsilon^{2}\ln\frac{1}{\varepsilon}$ for $m^{*}=4$ and depends $C^{1}$-continuously in $p_{t}^{M,k}$, $H^{M,k}_{j}$ and its derivatives up to $M+2$ order.
@@ 4.1 Comparison between $\varepsilon_{t}^{(1),k}$ and $\varepsilon_{t}^{(2),k}$
$\Lambda_{I,N}(\rho^{k})$ $=\sum_{i=1}^{n}\sum_{t=1}^{N}D_{it}^{k,(1)}\frac{e^{H_{it}(p^{(1),k})}}{e^{H_{i1}(p^{(1),k})}}(\varepsilon_{1}^{(1),k})^{m_{i}^{k}-2}+T_{M,k}^{(1),\Lambda}+O((\varepsilon_{t}^{(2),k})^{M(m^{*}-2){-\delta}})$
$=\sum_{i=1}^{n}\sum_{t=1}^{N}D_{it}^{k,(2)}\frac{e^{H_{it}(p^{(2),k})}}{e^{H_{i1}(p^{(2),k})}}(\varepsilon_{1}^{(2),k})^{m_{i}^{k}-2}+T_{M,k}^{(2),\Lambda}+O((\varepsilon_{t}^{(2),k})^{M(m^{*}-2){-\delta}}).$ (4.1)
Here, the terms $D_{it}^{k,(1)}$, $D_{it}^{k,(2)}$, $T_{M,k}^{(1),\Lambda}$, $T_{M,k}^{(2),\Lambda}$ and are defined as in Theorem 3.2.
@@
Recall that due to the notions in Theorem 3.2, we get
$T_{M,k}^{(l),\Lambda}=\frac{1}{2\pi N}\sum_{t=1}^{N}\sum_{i=1}^{n}(2-m_{t,i}^{k})\big(T_{M,k,i,t}^{(l),in}+T_{M,k,i}^{(l),\rho,out}\big)$
for $l=1,2$. Notice that the quantities $T_{M,k,i,t}^{in}$ and $T_{M,k,i}^{\rho,out}$ depend on $p^{(l),k}_{t}$ and $\varepsilon_{t}^{(l),k}$ for $l=1,2$. Then, a direct computation implies that
$|T_{M,k,i,t}^{(1),in}-T_{M,k,i,t}^{(2),in}|\leq A+B.$
@@
$|T_{M,k,i,t}^{(1),in}-T_{M,k,i,t}^{(2),in}|=O\Bigg(\varepsilon^{m^{*}-1}\Bigg|p_{t}^{(1),k}-\frac{\varepsilon_{t}^{(2),k}}{\varepsilon_{t}^{(1),k}}p_{t}^{(2),k}\Bigg|\Bigg).$
For terms $T_{M,k,i}^{(1),\rho,out}$ and $T_{M,k,i}^{(2),\rho,out}$, we get a similar result.
$|T_{M,k,i}^{(1),\rho,out}-T_{M,k,i}^{(2),\rho,out}|=O\Bigg(\varepsilon^{m^{*}-1}\Bigg|p_{t}^{(1),k}-\frac{\varepsilon_{t}^{(2),k}}{\varepsilon_{t}^{(1),k}}p_{t}^{(2),k}\Bigg|\Bigg).$
@@ 4.2 Comparison between $p_{t}^{(1),k}$ and $p_{t}^{(2),k}$
$\int_{B_{\tau_{k}^{M}/\varepsilon_{t}^{(1),k}}(p_{t}^{(1),k})}\partial_{1}H_{i}^{M,k}(y)e^{V_{t,i}^{(1),k}}=-T^{(1),Po}_{M,k}+O(\varepsilon^{(M+1)(m^{*}-2)-\delta})$
@@
$\int_{B_{\tau_{k}^{M}/\varepsilon_{t}^{(2),k})}(p_{t}^{(2),k})}\partial_{1}H_{i}^{M,k}(y)e^{V_{t,i}^{(2),k}}=-T^{(2),Po}_{M,k}+O(\varepsilon^{(M+1)(m^{*}-2)-\delta}).$
@@
$-T^{(1),Po}_{M,k}+T^{(2),Po}_{M,k}=O(\varepsilon^{m^{*}-2})\Bigg(\Bigg|1-\frac{\varepsilon_{t}^{(1),k}}{\varepsilon_{t}^{(2),k}}\Bigg|+|\varepsilon_{t}^{(1),k}p_{t}^{(1),k}-\varepsilon_{t}^{(2),k}p_{t}^{(2),k}|\Bigg).$ (4.3)
@@ 5.3 Higher order expansions via induction
Here, we use $\overline{T^{\mbox{in}}_{i,k,2}}(x)$ and $\overline{T^{\mbox{out}}_{i,k,2}}$ to denote the higher order expansions.
@@ A.3 A remark on a non-vanishing blowup result
In [12], the authors glue a blowup solution under the assumptions $(\mathcal{H}_{1})$, (2.1), (2.2) and (3.13). The assumption (2.1) does not require